Home‎ > ‎Reports‎ > ‎crew110b‎ > ‎

day02


Crew #110B Log Book for January 8, 2012

Photos of the day





Commander's Check-In Report

Date and time: 8 January 2012, 17:15

Written by: April Andreas, Crew 110B, Crew Commander

Crew Physical Status: In good spirits.

Time Departed/Returned from EVA:

We went on two EVAs. Each EVA had the same goal: in-sim orientation and GPS navigation practice and testing of our communication equipment. EVA 1 departed 10 am, returned 12:10 . EVA 2 departed 13:52 , returned 16:09 pm.

Brief Narrative of Field Mission Results:
Each team visited the same sites and did minor collector of gypsum, etc.

EVA Data/Interpretations:
None to report, individual EVA leads may have more detail.

Engineering/Hab Maintenance:
We request permission to adjust the idle speed on ATV 2, since it is too low, and it does not idle on its own. We do need more water outside. We were concerned that the diesel gauge may not be reading correctly, since it's been reading 0.75 for about a week, according to Crew 110A.

Report Transmission Schedule:
As completed.

Plans for Tomorrow:
Tomorrow we're going to take swabs in the hab to test for a variety of coliforms and staphlocci. In the afternoon half our crew will retrace our steps of today's EVA to create a "guided tour" EVA for our students. The other half will stay at the hab to document the microbiology work and create a plan for using the telescope.

Inventory:
n/a

Support Requested:
* We need a new shipment of fresh water. It's at less than half. We had to use a lot of water getting the lab clean yesterday, washing handtowels, etc., and that depleted the supply considerably. It's at 9.5 inches.
* We request permission to adjust the idle speed on ATV 2.

Miscellaneous:
n/a

Commander's Report

Date and time: 8 January 12, 17:45

Written by: April Andreas, Crew 110B, Commander

Text:

Today was a fantastic day with the crew. We completed our first two EVAs, for which the primary purpose was EVA and ATV orientation, in-sim. Even though we had about an hour allocated to figure out how to get the suits on for the first time, it actually took about an hour and twenty minutes! We used athletic tape to secure the headsets for our two-way radios, but were disappointed that CAPCOM couldn't stay in communication with the EVA crew for a longer range.

We are trading off the different roles and responsibilities of the crew, so Liz and Nancy have been following up with the engineering checklists. Elaine and Larry are figuring out how to prepare the meal tonight -- we are really looking forward to the brownies!

Tomorrow we'll start doing our first lab experiment to test for a variety of coliforms and staphlocci. In the afternoon half our crew will retrace our steps of today's EVA to create a "guided tour" EVA for our students. The other half will stay at the hab to document the microbiology work and create a plan for using the telescope.


EVA 1 Report

Date and time: 8 January 2012

Written by: Elaine Fagner, Crew #110b; Crew Position: Geologist/HSO

Time: 10:00 - 12:10

Crew Members: Vernan Kramer; Larry Benton, Elaine Fagner

Site Location: North of Hab with multiple stops.

Transit Mode: ATV

Objectives: Initial reconnaissance of the territory and region to help assist in specific scientific problems and projects.

Results: Team observed specific geologic depositional environments and collected specimens (rocks, minerals) for study. We left later than our assigned time due to radio problems with Vernan's equipment. We pressurized twice and Vernan had to go back in another time to fix it again. We arrived 10 minutes later due to Vernan's watch not being syncd with the Hab. Team radioed the Hab for time check and our communications were not received. They were able to reach us by 12:00 and we were already in transit back to the Hab.

Lessons Learned: It would be helpful to have driving hand signals that are used in case one of the radios is not operating properly. Elaine's radio kept going into constant Voice Activated Mode so communications were difficult between members. Having a mirror for the EVA lead would be helpful to see the status of other team members.

EVA 2 Report

Date and time: 8 January 2012

Written by:  Liz Mitchell , Crew #110b; Crew Position:  Biologist/HSO

Time:  13:52 - 16:10

Crew Members:  Liz Mitchell, Nancy Ray-Mitchell, April Andreas

Site Location:  North of Hab with multiple stops.

Transit Mode:  ATV

Objectives:  Initial reconnaissance of the territory and region to help assist in specific scientific problems and projects.

Results: Team took pictures of several sites along the road, for PR purposes to promote our program on our return. We also wanted to get team members comfortable with the ATVs and using the GPS. We left later than our assigned time due to the late arrival of the previous team. We had to have a debriefing with that team.  We also had to re-fill ATVs with gas. During the return trip, Liz's radio became disconnected from her ear. She was unable to replace it. We pressurized both leaving and returning. After about 1 mile we lost communications were not received. We resumed communication at around the same location at which we lost communication upon return.

Lessons Learned: We learned how to operate the GPS and ATVs properly. It would be helpful to have driving hand signals that are used in case one of the radios is not operating properly. Having a mirror for the EVA lead would be helpful to see the status of other team members. 


Engineering Report



Journalist Report

Date: 8 January 2012
Written by: Nancy Ray-Mitchell

The team ate breakfast together and discussed plans for the day. We went over GPS coordinates and the planned routes for today's EVAs. It took about an hour and 20 minutes for Liz, April and Nancy to help Vernon, Larry, and Elaine get ready for the first EVA.

The first group departed about 30 minutes behind schedule due to complications with our radio equipment. They had to pressurize twice due to a microphone issue with one of the crew members. The major destinations today were gypsum fields.

The CAPCOM and remaining crew communicated with the EVA crew as long as the radios allowed. The remaining crew practiced on the GPS, took pictures, cleaned the kitchen, and made lunch. The majority of the time was spent discussing feedforward control for bringing students to MDRS in the future.

The remaining crew helped the first crew get out of the spacesuits, replaced and charged all equipment, tried on spacesuits; and gassed the ATVs. Lunch was spent discussing what should be done differently for the second group going out. Radios needed to be on the right hand for easier communication; less clothing should be worn; more secured ski caps; and in general, we need to allow more time for EVAs.

The second crew (April, Liz, and Nancy) was able to get ready a little quicker because of lessons learned from the first group. After one missed turned, the second group was able to follow the GPS coordinates exactly. Upon returning, members of the second crew were individually assigned to one person in the first crew. It made getting undressed much easier and quicker.

As our two cooks for the night, Larry and Elaine, made our defined meal, the rest of us wrote reports, downloaded pictures, did engineering duties, and worked on battery packs. We ate together, discussed our great day, and discussed our plans for tomorrow in more detail.


Chef's Report

Date: 8 January 2012
Written by Elaine Fagner and Larry Benton of Crew 110B

1a) Today was
- a cooking day (designated meal)

1b) Was today a special day celebrated at the main meal with special food or activity? If so, what was special about today and what food(s) were prepared to mark the celebration? No.

2) List the foods served at today's main meal (usually the evening meal), giving full names of each.

Supper was a designated meal consisting of brown rice, rehydrated vegan chili, rehydrated broccoli and beef, rehydrated beef, and brownies.


3) List any main meal foods not finished by the crew (leftovers stored for later use or discarded) Chili, broccoli and beef, brownies.

4) Recipes for anything more complicated than following package directions exactly, or rehydrating.

Added 1 heaping teaspoon of McCormick Beef Base to the rehydrated beef to enhance the flavor.

5) Lessons learned

Cooking the brown rice took an extremely long time. Cooking of the rice was finished using the pressure cooker that was delivered with the food for Crew 110B. Some difficulty was experienced getting the lid on and off of the pressure cooker due to a locking apparatus. Further examination of the pressure cooker will be performed to determine correct operating procedures.

6) Comments/questions for Kim and Jean, the food study investigators

Brownies were a big hit, and the brown rice and rehydrated beef was well received. The broccoli and rice dish and the chili was not as well liked.